OpenStudy (anonymous):

MEDAL AND FAN Which of the following is a solution to the system of linear equations below? 2x=-3y-11 2y=6x-22

OpenStudy (anonymous):

What are the answer choices

OpenStudy (anonymous):

(2, 5) B. (–2, 5) C. (2, –5) D. (–2, –5)

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

I got (2,-5)

OpenStudy (anonymous):

2y = 6x - 22 y = 3x - 11 2x + 3•(3x-11) = -11 11x = 22 x=2 y = 3(2)11 = -5

OpenStudy (anonymous):

Solve the system by substitution. y = 2x 3x + 2y = 21 A. x = 3, y = 6 B. x = 6, y = 3 C. x = 3, y = –6 D. x = –6, y = 3

OpenStudy (anonymous):

Okay You plug the first equation into the second one which gives you 3x+2(2x)=21 3x+4x=21 7x=21 x=3 So now plug 3 into the first equation to get y y=2(3) y=6 (3,6)